A 10-year U.S. Treasury bond with a face value of $1,000 pays a coupon of 5.5% (2.75% of face value every six months). The reported yield to maturity is 5.2% (a six-month discount rate of 5.2/2 = 2.6%). (Do not round intermediate calculations. Round your answers to 2 decimal places.)

a. What is the present value of the bond?

Present value            $

b. If the yield to maturity changes to 1%, what will be the present value?

Present value            $

c. If the yield to maturity changes to 8%, what will be the present value?

Present value            $

d. If the yield to maturity changes to 15%, what will be the present value?

Present value            $
Answer #1
a) we have to use financial calculator to sovle this problem
Put in calculator
FV 1,000
PMT =1000*2.75% 27.5
I =5.2%/2 2.60%
N =10*2 20
compute PV ($1,023.16)
Therefore present value = $1,023.16
b) we have to use financial calculator to sovle this problem
Put in calculator
FV 1,000
PMT =1000*2.75% 27.5
I 1%/2 0.50%
N =10*2 20
compute PV ($1,427.22)
Therefore present value = $1,427.22
c) we have to use financial calculator to sovle this problem
Put in calculator
FV 1,000
PMT =1000*2.75% 27.5
I =8%/2 4.00%
N =10*2 20
compute PV ($830.12)
Therefore present value = $830.12
1) we have to use financial calculator to sovle this problem
Put in calculator
FV 1,000
PMT =1000*2.75% 27.5
I =15%/2 7.50%
N =10*2 20
compute PV ($515.76)
Therefore present value = $515.76